OpenStudy (anonymous):

I need help with solving this radical expression. I have been stuck on it for two days! sqrt(8x-23)=4-x

OpenStudy (anonymous):

The answer is 3, however I don't know how they got it.

OpenStudy (ivettef365):

what is 3

OpenStudy (austinl):

\[\sqrt{8x-23}=4-x\] \[(\sqrt{8x-23})^2=(4-x)^2\] That gets rid of the square root, and then you need to FOIL the right side. Can you do that for me?

OpenStudy (austinl):

(4-x)(4-x)=?

OpenStudy (anonymous):

8x-23=16+x^2

OpenStudy (anonymous):

or, 8x-23= 16-4x-4x=x^2

OpenStudy (austinl):

\[8x-23=x^2-8x+16\] \[-x^2+16x-39=0\] Then you should be able to factor and solve for the answer. Or quadratic equation.

OpenStudy (phi):

** or, 8x-23= 16-4x-4x=x^2 *** I assume that last = is a typo ? It should be + 8x-23= 16-4x-4x+x^2 on the right side, -4x - 4x combines to make -8x 8x-23= 16-8x+x^2 switch sides (because that is how I am used to looking at these equations) (and put in standard order) x^2 - 8x +16 = 8x -23 add -8x +23 to both sides so that everything is on the left side x^2 -8x -8x +16 +23 = 8x-23-8x +23 simplify x^2 -16x +39 = 0 can we factor this ? list pairs of factor so 39 1,39 3,13 look for a pair that add to 16. the 2nd pair work. make both - because we want -16 (x-3)(x-13)= 0 can you finish ? remember to check the solution in the original equation, because the square root could cause troubles if you take the square root of a negative number

OpenStudy (phi):

also, for this type equation sqrt(8x-23)=4-x you only take the + square root (the principal root) that means you exclude any x values that make the right side negative ... because the left side is assumed to always be positive.

OpenStudy (anonymous):

i did have a typo on the last one. I think I must have completely failed the quadratic equations section. that is where I lost it.

OpenStudy (anonymous):

so since one of the blocks needs to be zero, (x-3)or (x-13) to equal zero I find that 3 is the only one that works in the original equation. sqrt of 8*3-23=4-3 1=1 the other solution of 13 does not work since 9 does not equal -9 THANKS!